The area of a given hexagon is equal to the area of an equilateral triangle whose perimeter is 36in, find the length of the side of a regular hexagon
Answer by mananth(16946)

equilater triangle
perimeter = 36
so each side = 12in
..
height of equilateral triangle = (sqrt3)s /2 s= side of triangle
h= 6*sqrt3 / 2
h= 3sqrt3
= 5.2 in
..
area of triangle = 1/ * base * height
= 1/2 *6 *5.2
=3*5.2
=15.6 sq.in
..
so area of hexagon = 15.6
area of hexagon = 3sqrt(3)/2 * s^2 s= side of polygon
=36*3*sqrt(3) / 2
= 93.53 sq. inches